OpenStudy (anonymous):

Factor and simplify. Express the answer as a fraction without negative exponents. x(x-1)^1/2 + 2(x-1)^1/2 are you supposed to do this: (x+2) sqrt(x-1) / sqrt(x-1) ??? but then what?

OpenStudy (dumbcow):

where did the sqrt come from ??

OpenStudy (dumbcow):

oh nevermind, i misread the expression :)

OpenStudy (dumbcow):

ok you dont need the sqrt(x-1) on bottom

OpenStudy (anonymous):

(x-1)^1/2(X+2)

OpenStudy (anonymous):

oh crap i meant, x(x-1)^-1/2 + 2(x-1)^1/2, haha sorry >.<

OpenStudy (dumbcow):

\[\frac{x}{\sqrt{x-1}}+2\sqrt{x-1} = \frac{x}{\sqrt{x-1}}+(2\sqrt{x-1}*\frac{\sqrt{x-1}}{\sqrt{x-1}}) = \frac{x+2(x-1)}{\sqrt{x-1}}\]

OpenStudy (dumbcow):

\[=\frac{3x-2}{\sqrt{x-1}} = \frac{3x-2}{\sqrt{x-1}}*\frac{\sqrt{x-1}}{\sqrt{x-1}} = \frac{(3x-2)\sqrt{x-1}}{x-1}\]

OpenStudy (anonymous):

haha wow thanks so much!

OpenStudy (dumbcow):

i was assuming they won't like radical in the denominator your welcome